x^4 - 10x^2 + 9 / x^3 + 4x^2 + 3x
Answer by Linz(9)

For x^4-10x^2+9 we need to factor it to (x^2-9)(x^2-1)
Then factor x^3+4x^2+3x
Since there is an x in each term we can factor it out
x(x^2+4x+3)
We need to factor both top and bottom more.
(x^2-9)(x^2-1)= (x+3)(x-3)and(x+1)(x-1)
x(x^2+4x+3)= x(x+3)(x+1)
then you can cancel out terms that are the same in top and bottom.
We cancel out (x+3) and (x+1) in the top and bottom to get an answer of:
(x-3)(x-1)/x
Hope this helps!